OpenStudy (kamille):

How many even numbers can be made from number 2567 if digits can't repeat. How I think: One digit number: 2 Two digits number: 2*2=4 how to find three and four digit numbers?

OpenStudy (anonymous):

if the last digit is 2 then there are \(3!=3\times 2=6\) choices for the other three

OpenStudy (anonymous):

and if the last digit is 6, then there are still \(3!=6\) choices so the total is \(6+6=12\)

OpenStudy (kamille):

well, i did it like it, but there numbers with 4 digits, right? And the answer is 32, so I think something is incorrect

OpenStudy (anonymous):

oh sorry, i thought you had to use all the digits

OpenStudy (kamille):

oh, actually i dont know. the problem sounds like this: "How many even numbers can be made from number digits 2567, if digits cant repeat". Do I need to use all numbers? Cant it be any one digit, two digit numbers?

OpenStudy (kamille):

@satellite73 are you here?

OpenStudy (kamille):

@hartnn

OpenStudy (kamille):

can you help me?

hartnn (hartnn):

even i get 12 even numbers using same logic as of satellite.

OpenStudy (kamille):

Oh, i think it can be: One digit numbers: 2, 6 (2 numbers), two digit numbers (12;16;26;56;52, etc) - 8 numbers, three digit numbers (12 of them). Isnt it correct?

hartnn (hartnn):

oh, so not necessarily 4 digit number. you need to add number of 1 digits nos. + number of 2 digits nos. +number of 3 digits nos. +number of 4 digits nos. and 12 is the number of 4 digits number.

OpenStudy (kamille):

Can you show me step by step?

hartnn (hartnn):

One digit number: 2 Two digits number: 3C1*2=6 3 digit number = 3C2 * 2! = 6*2 = 12 4 digit number = 3! *2! = 12 2+6+12+12 = 32

OpenStudy (kamille):

thanks:)

hartnn (hartnn):

only 6 two digit number 12;16;26;56;52;62

hartnn (hartnn):

welcome ^_^